### v1

# This Candidate is [MASK]

Abstract
I implement a prompt-based learning strategy to extract measures of sentiment and other features from confidential reference letters. I show that the contents of reference letters is clearly reflected in the performance of job market candidates in the Economics academic job market. In contrast, applying traditional “bag-of-words” approaches produces measures of sentiment that, while positively correlated to my LLM-based measure, are not predictive of job market outcomes. Using a random forest, I show that both letter quality and length are predictive of success in the job market. Letters authored by advisers appear to be as important as those written by other referees.
JEL classification:
Key words: Large language models; text data; sentiment analysis.
1 Introduction
Accessing job seekers’ social networks through references is an important way employers can expedite hiring and reduce search costs. Hiring in real-world labor markets is a costly process characterized by considerable uncertainty regarding the aptitude, skill, and motivation of job applicants. Firms can advertise job vacancies widely and then proceed to a meticulous process of selection through interviewing and testing candidates. Alternatively, firms can economize on search costs by eliciting recommendations from trusted sources. References can come from current and former employees, or from experts in the field. [^1] There are, of course, many other possible strategies to address search frictions, such as the use of external recruiters or “raiding” the employees of competitors (see Oyer and Schaefer, 2011).
The importance of references in the labor market has been known to economists and sociologist for several decades (Rees, 1966, Granovetter, 1973). Widely cited studies put the prevalence of references as a way of filling positions well above 50%. [^2] See Ioannides and Loury (2004) and Topa (2011) for excellent reviews of the literature.
In this paper, I focus on the use of reference letters in the context of the job market for academic economists. Using an administrative data set covering all applications to a top economics department in Russia and a new prompt-based strategy to extract sentiment scores from reference letters with a large language model, I provide direct evidence that the contents of letters of recommendation are predictive of job market outcomes.
My LLM-based method has multiple advantages relative to existing methods for sentiment analysis in economics. Language models are trained to understand context, and are not focused on the presence or absence of certain words — the essence of widely applied “bag-of-words” methods. Second, by using different prompts it is possible to extract information on observable features — like the sex of the candidate — which provides a test of the model’s predictive power. Third, because language models output probabilities, the resulting sentiment scores have a clear interpretation and measurement units. Finally, prompting may be used to extract other predictive features from text, not just sentiment.
The economics job market is an ideal setting to understand how references are used as a way to screen applicants. Because all job market candidates have similar formal credentials, the use of education as a signalling device is implicitly controlled for. As a way to further control for the level of training, I focus only on first time job market participants. In other words, I only look at job market candidates who are close to finishing their PhD program and have not held a permanent academic position before.
The job market for economists is somewhat unique in that all open positions are widely advertised. With few exceptions, all openings are posted in a central database accessible to all market participants. Moreover, the application process itself has become highly standardized. Candidates are expected to submit a CV, a research sample (“job market paper”), and few other materials. [^3] For example, some colleges with teaching-focused positions might require a “teaching philosophy” statement and the submission of student evaluations. The extensive use of electronic applications means that applying is almost costless. In summary, the specific characteristics of the market mean that references are not being used to disseminate information about job market opportunities. [^4] For a model of the role of social networks as a manner of obtaining information about job opportunities, see Calvó-Armengol and Jackson (2004). Instead, reference letters are used as a way to screen applicants. [^5] There are alternative explanations for the use of references that do not focus on either search costs or screening. Kugler (2003) and Heath (2018) consider the possibility that firms use references as a way to lower monitoring costs and improve labor discipline. Kramarz and Skans (2014) and Wang (2013) provide evidence that in certain contexts family relations are important in finding jobs. Neither of these explanations seem particularly relevant to the economics job market.
Aside from isolating the screening channel, the economics job market is unique in that the use of references has been formalized. Indeed, job market candidates who wish to apply for a position are customarily instructed to have two or more confidential letters of reference submitted directly to the hiring committee. Because candidates typically submit the same letters to all positions they apply to, I have precise information on who wrote letters for each job market candidate in my sample. In contrast, many prior studies do not observe recommendations directly and are therefore forced to use some kind of proxy. [^6] For example, Dustmann et al. (2016) define networks in terms of ethnicity, while Bayer et al. (2008) and Hellerstein et al. (2011) use the neighborhood of residence as proxy.
The issue of under what conditions indirect signaling via letters of recommendation can function efficiently has been a matter of concern in the theoretical literature. Because referees compete against each other, we can expect them to make excessive claims about the workers they are recommending in order to have them placed ahead of the candidates being recommended by a rival referee (Saloner, 1985). [^7] Other models in which referrals serve as a screening mechanism to reduce the information asymmetry inherent to the hiring process are Montgomery (1991) and Simon and Warner (1992). Moreover, referees’ motivation for referring applicants and providing truthful recommendations is not well understood. Câmara et al. (2023) model the reputation game facing referees explicitly. Remarkably, they find that referees may lie in equilibrium, and that the extent of the lying depends on the intricate interplay between preferences, reputation and lying costs. [^8] Another paper that investigates referees’ incentives theoretically is Ekinci (2016). Carmichael (1988) argues that the reason why universities have instituted a tenure system in academic departments is precisely to ensure that faculty members are properly incentivized to hire the best job market candidates.
In line with these considerations, the extent to which the strength of an applicant’s references correlate with good labor market outcomes is an empirical question without an obvious a priori answer. If hiring committees learn to distrust the contents of reference letters because of generalized “recommendation inflation,” then the correlation might be statistically indistinguishable from zero. On the other hand, if reputational and other costs are severe enough to align the incentives of referees with those of hiring committees, then we can expect to find that good references more often than not translate into success in the job market.
There are several empirical studies of how references affect labor market outcomes. Using a long panel data set covering all hiring activity of a single US corporation, Brown et al. (2016) found that referred candidates are more likely to be hired than non-referred candidates. Burks et al. (2015) assembled personnel data from nine large firms in the call center, trucking, and high-tech industries. Consistent with Brown et al. (2016), they found that referred applicants are substantially more likely to be hired, and that, conditional on receiving an offer, they are more likely to accept it.
There is also some experimental evidence on the value of reference letters. Pallais (2014) conducted a field experiment on an online marketplace in which workers bid to be hired for short term employment. One of her treatment groups received detailed feedback after completing a data-entry task, while another group received only a one-to-five rating. She found that, in the two months after treatment, both treatment groups benefitted from treatment relative to a control group that was not hired. Treated workers worked more jobs and had higher earnings. In addition, making detailed positive feedback on past performance publicly available further increased their employment prospects. Abel et al. (2020) conducted an audit study where they randomly appended a reference letter to some applications submitted on behalf of job seekers to vacancies in South Africa. They found that, for the same applicant, attaching a letter increased the probability of receiving a response and the rate of interview requests by about 60 percent.
Two papers study the use of reference letters in the economics job market (Eberhardt et al., 2023, Baltrunaite et al., 2024). Their data is similar to mine in that they collect administrative hiring data from institutions hiring in the international economics job market. [^9] Eberhardt et al. (2023) study a sample of applicants to a top UK economics department, while Baltrunaite et al. (2024) look at a premier economics department and a top policy institution in Italy. While these studies do link reference letter features to job market outcomes, their focus lies more on gender differences in the linguistic contents of reference letters and their consequences.
This paper also makes a methodological contribution to the broader literature that makes use of text data in economics and finance. I show that a prompt-based strategy can take advantage of a LLM’s pre-training to extract sentiment and other scores from the text. This new strategy presents several distinct advantages over other methodologies that have been deployed in economics and finance.
In summary, I find that traditional dictionary-based methods are unable to extract information from reference letters that is predictive of job market outcomes. Prompt-based sentiment extraction, in contrast, produces sentiment scores that are strongly correlated with the probability of obtaining a permanent research position. The relationship is even stronger when predicting employment at high rank institutions.
Until fairly recently, feasible strategies to take advantage of text data were rather limited. In the next section, I provide a brief review of the existing methods and a formal description of my prompt-based strategy. I then move to a description of the data, including a descriptive analysis of my prompt-based sentiment score. Section 4 presents the main results and section 5 concludes.
2 Methods: Sentiment Analysis and Prompt-based Learning
In this paper, I apply a prompt-based methodology to extract sentiment and other scores from reference letters. This approach has several advantages over existing methods. In this section, I first provide a brief description of the lexical approach, which has been the most commonly applied method to assign a sentiment score to text in the economics and finance literature. I then describe the prompt-based methodology for sentiment extraction.
2.1 Dictionary Methods
Text analysis in economics has a long history. [^10] The classic Gentzkow et al. (2019) and the more recent Ash and Hansen (2023) provide great reviews of the different techniques that have been deployed to handle text in economics and finance research. In particular, several papers have studied the problem of how to assign a sentiment score to text.
Shapiro et al. (2022) define sentiment as “a measure of the speaker’s tone, attitude, or evaluation of a topic, independent of the topic’s own sentiment orientation.” The main approach to sentiment analysis in economics and finance has been the lexical methodology, which relies on a preset dictionary of words, with each word assigned a sentiment score. [^11] Typically, words in the dictionary have a score of 1 if they are deemed to express positive sentiment, -1 if negative, and 0 for neutral words. But more complex approaches exist. Within this framework, the sentiment of a document is some simple aggregate of the scores of its constituent terms. The most common approach is to use the polarity score:
[…]
where $S^{i}_{d}$ represents the sentiment score assigned by the lexicon to term $i$ of document $d$. For the simple case of ±$\pm 1$ scores, this simplifies to
$\frac{N^{+}_{d}-N^{-}_{d}}{N^{+}_{d}+N^{-}_{d}},$
where $N^{+}_{d}$ and $N^{-}_{d}$ are, respectively, the counts of positive and negative terms. The methodology is often referred to as “bag of words” because the order in which the terms appear does not affect the resulting sentiment score.
The literature has devoted considerable attention to the quality and domain-specificity of the employed lexicons. For instance, Loughran and McDonald (2011) showed that the Harvard General Inquirer (HGI, a commonly used dictionary) does not appropriately evaluate the sentiment in financial/economic text. Instead, they developed their own lexicon of terms for economics and finance, which to this day remains an important reference point.
Another attempt at improving on the lexicon approach adds a series of heuristic rules to take into account common linguistic patterns that affect sentiment, such as negation and emphasis. A prominent example is the Vader analyzer (Hutto and Gilbert, 2014), which is designed to deal with short texts such as tweets and online reviews. In this paper, I also use SenticNet, which uses a “concept-based” lexical approach. [^12] See sentic.net for details.
Generally speaking, even the most sophisticated lexical approach does not take into account the context in which words occur. It is not difficult to construct examples where several seemingly positive terms are used in a phrase in order to convey a strong negative opinion. Sarcasm, understatements, backhand comments, etc. are undetectable to lexicons.
In addition to this limitation, bag-of-words methods also suffer from the fact that the resulting sentiment score does not have an easy interpretation or unit of measurement. While researchers customarily compare scores resulting from alternative lexicons, it is not in general possible to evaluate the relative accuracy of the different sentiment scores.
A final limitation of the lexical approach is that, if the method is to be used to generate a measure other than a sentiment score, a completely new dictionary must be produced. For example, Schmader et al. (2007) created a set of dictionaries to summarize different semantic aspects of reference letters written for chemists and biochemists. Specifically, they developed sets of terms associated with the following language categories: grindstone traits, ability traits, research terms, teaching and citizenship terms, and standout adjectives. Eberhardt et al. (2023) applied a similar strategy to letters of recommendation written on behalf of job market candidates applying to a top UK economics department.
[…]
2.2 Prompt-based Learning
In recent years, there has been a paradigm shift in natural language processing (NLP), inspired by the development and fast popularization of large pre-trained transformer-based language models (PLMs). In a nutshell, the paradigm shift involved moving from a pre-train then fine-tune strategy to the prompt-based learning approach. [^14] For the sake of brevity, I limit the description of the paradigm shift and of prompt-based learning to the bare minimum. See Sun et al. (2022), Min et al. (2023) and Liu et al. (2023) for comprehensive reviews. Bu et al. (2024) focus on sentiment analysis via prompt learning.
The idea of pre-training is to first train one large model on a “fundamental” pre-training task and only later adapt (“fine-tune”) it to a multiplicity of specialized tasks in a second step. In the context of NLP, the pre-training task has always been some kind of language modeling, which involves either unidirectionally predicting the next word in a sequence of text, or a fill-in-the-blank style task in which the model must predict a masked word in a sentence. [^15] The ULMFiT model (Howard and Ruder, 2018) is widely credited with popularizing pre-training and fine-tuning for NLP tasks. That model pre-trained by predicting the next token in a sequence. The great advantage of pre-training is that the model can exploit the massive quantities of text available from the web and other sources, largely obviating the extremely expensive labeling required in classical supervised learning.
The pre-training technique has been instrumental in the development of large language models following the development of the transformer architecture (Vaswani et al., 2017). The transformer’s attention mechanism allows the processing of text sequences without the computationally expensive recurrent processing required in LSTMs and the other recurrent neural network layers that were the previous state of the art. [^16] See Zhao et al. (2023) for a detailed review of the recent advances of LLMs. Importantly, soon after the first transformers were rolled out, it became apparent to researchers that the models had capabilities that far exceeded simple language generation. The newly discovered “emergent properties” of LLMs range from text classification to translation tasks (see Brown et al., 2020).
Instead of fine-tuning a PLM to downstream tasks, these tasks can be reformulated to look more like those solved during the model’s pre-training. The key insight is that a textual prompt can be devised to elicit the desired prediction. For example, in order to recognize the emotion of the tweet “I missed the bus today!”, we may append the prompt “I felt so [MASK]” and ask the LLM to fill the blank with an emotion-bearing word. [^17] The same prompting strategy can be applied to multiple tasks. To elicit a translation, one might use the prompt: “English: I missed the bus today! French: …” and use a generative LLM to generate the translation.
The great advantage of reformulating the problem this way is that, once the prompting strategy has been set, one may proceed to making predictions without the need of any further training. Specifically, prompting spares us from the often unsurmountable task of collecting a large labeled data set. Moreover, in many cases —including assigning sentiment to confidential reference letters— the difficulty inherent in the labeling process itself may become an obstacle.
To give a formal statement of these points, consider first a supervised learning NLP task that takes a text $\boldsymbol{x}$ and predicts output based on a model $\mathrm{P}\left(y\mid\boldsymbol{x};\boldsymbol{\theta}\right)$. For a text classification task, given an input text $\boldsymbol{x}$ the model predicts a label $y$ from a fixed label set $\mathcal{Y}$. For example, a sentiment analysis task might take an input such as “I loved this movie!” and predict a label $y=+$ out of the label set $\mathcal{Y}=\{+;\sim;-\}$.
In this terms, the challenge of supervised learning is that in order to find optimal weights $\boldsymbol{\theta}$, we need lots of supervised data $\left(\boldsymbol{x}_{i},y_{i}\right),i=1,\ldots,N$ where the $y_{i}$ are the labels (a.k.a. “ground truth”) for a number of training examples.
Prompt-based learning takes advantage of a pre-trained model $\mathrm{P}\left(\boldsymbol{x};\boldsymbol{\gamma}\right)$ of text $\boldsymbol{x}$ itself. The derivation of predicted probabilities requires three steps:
Step 1.
The first step involves the specification of a prompting function $f_{prompt}(\cdot)$. The prompting function takes text $\boldsymbol{x}$ as an input and produces a prompt: $\boldsymbol{x}^{\prime}=f_{prompt}(\boldsymbol{x})$. The easiest way to specify $f_{prompt}(\cdot)$ is to use a template ($\mathcal{T}$), i.e. a text string that contains two slots: an input slot [X] for the input $\boldsymbol{x}$ and an answer slot [MASK] for an intermediate answer text $\boldsymbol{m}$.
For example, for the task of assigning sentiment to movie reviews, a suitable template would be: “[X] Overall, it was a [MASK] movie.” The resulting prompt would be
[…]
Step 2.
[…]
Step 3.
The third and final step is simply to collect predicted probabilities for each label using the PLM. Let the function $f_{fill}(\boldsymbol{x}^{\prime},\boldsymbol{m})$ fill-in the location [MASK] in prompt $\boldsymbol{x}^{\prime}$ with the permissible answer $\boldsymbol{m}$. The predicted probabilities for each label are thus defined:
$\mathrm{P}(y\mid\boldsymbol{x})=$ $\sum_{\mathbf{m}\in\mathcal{M}}\mathrm{I}\left[\mathcal{V}(\boldsymbol{m})=y\right]\cdot\mathrm{P}\left(f_{fill}(f_{prompt}(\boldsymbol{x}),\boldsymbol{m});\boldsymbol{\gamma}\right),\quad y\in\mathcal{Y}$ (1)
where $\mathrm{I}[\cdot]$ is the indicator function. Intuitively, the prompt-based probabilities are a measure of how likely the PLM judges each of the filled prompts. In particular, the predicted probability of assigning positive sentiment to the input text is simply the sum of the probabilities — based on the LLM’s pre-training — assigned to each of the positive tokens in the verbalizer. This intuitive interpretation is one of the advantages of prompt learning. In addition, because sentiment scores are derived from the predicted probabilities, they have a natural unit of measurement.
The predicted probabilities can be used to predict a sentiment class $\hat{y}=\arg\max_{y\in\mathcal{Y}}\mathrm{P}(y\mid\boldsymbol{x})$. However, in sentiment analysis it is customary to use the text’s polarity [^18] Polarity is a sentiment measure that is robust to over-sensitivity to subjective or otherwise sentiment carrying tokens in the input text., defined as:
$\mathrm{polarity}(\boldsymbol{x})=$ $\mathrm{P}\left(y=+\mid\boldsymbol{x}\right)-\mathrm{P}\left(y=-\mid\boldsymbol{x}\right)$
2.3 Applying Prompt-learning to Reference Letters
In principle, any PLM capable of evaluating $\mathrm{P}\left(\boldsymbol{x};\boldsymbol{\gamma}\right)$ can be used for prompting. However, the model choice has important implications for the overall quality and precision of the sentiment scores, as well as for other somewhat more technical aspects.
  1. 1.
    Generative models such as those in the GPT and Llama families require a prefix prompt. In other words, the [MASK] term in the template must be the last token. [^19] Because of this restriction, the [MASK] term is often omitted altogether. The reason for this is that generative models are pre-trained using causal language modeling, i.e. predicting the next word in text sequences. In contrast, bidirectional models like Bert, Roberta, or Big Bird are pre-trained via masked language modeling, so the [MASK] term can be anywhere in the template. [^20] These are called cloze prompts.
  2. 2.
    LLMs operate through a pipeline that begins with the tokenization of the input text. Each model’s tokenizer is somewhat different. The main restriction here is that the terms in the permissible set $\mathcal{M}$ must be in the tokenizer’s vocabulary. [^21] One important issue to check is whether multiple spellings of the same word are in the vocabulary. For example, both “labor” and “labour” are in most LLMs vocabularies.
  3. 3.
    An important consideration is that the length of the input text (measured in tokens) must fit within the model’s context size, which varies from model to model.
In this paper, I use Big Bird (Zaheer et al., 2020), a bidirectional LLM pre-trained on Wikipedia, books, short stories, and news articles. Big Bird’s vocabulary contains over 50 thousand tokens, including almost every common-use English word. [^22] Big Bird’s tokenizer is based on the SentencePiece algorithm, so some tokens in the vocabulary are common word parts or punctuation.
Big Bird implements a special attention layer (named “sparse attention” by the Google team that developed it) with a context size of 4,096 tokens. The model can also be used in “full attention” mode, which reduces the context size to 1,024 tokens but increases performance. [^23] Because many documents in my data set exceed the 1,024 token limit, using the model in full attention mode requires applying a technique known as windowing. More details are provided below.
2.4 Sentiment Prompt and Verbalizer
To extract sentiment from reference letters, I use the following template/verbalizer combination:
$\mathcal{T}_{sentiment}:=$ “[X] In summary, this job market candidate is [MASK].”
$\mathcal{V}_{sentiment}:=$
    excellent, outstanding, exceptional, brilliant, great, superb, fantastic, terrific, amazing, phenomenal, awesome, stellar, wonderful, extraordinary, remarkable, impressive, incredible, fabulous, formidable, marvelous, perfect, fascinating.
    bad, terrible, disappointing, poor, awful, lousy, atrocious, pathetic, unsatisfactory, substandard, mediocre, common, inferior, average, ordinary, regular.
3 Hiring Data
The analysis in this paper is based on administrative data from my institution’s recruitment efforts during the 2013, and 2015–2021 job markets. [^24] I designate job market cycles by the year in which recruitment formally started and the job ads were posted. Unfortunately, data for the 2014 market was not archived. To protect confidentiality, I processed all data myself.
As already mentioned, I only consider job market candidates who are finishing their PhD programs and have not held a full-time academic position. In addition, only applications where the adviser’s letter of recommendation and at least one additional letter were submitted are kept. Overall, my data contains 1,998 letters from 603 job market candidates.
The pre-processing of the reference letters’ text was kept to a minimum. I removed headers, footers, page numbers, salutations, and some obvious mistakes from faulty conversions from pdf format. I also collected personal data from the submitted CVs, as well as personal websites and LinkedIn profiles. Specifically, I have data on the candidate’s sex, nationality, prior publications and time to PhD completion. I also collect the PhD granting institution’s location and ranking. [^25] To rank Economics departments, I use Tilburg University’s ranking, which is based on publications weighted by journal impact factor and covers the whole world (https://rankings.tilburguniversity.edu/).
Finally, I hand-collected the detailed job market outcome for each candidate in my data. The job market outcome corresponds to the job held by September of the following year.
3.1 Descriptive Statistics
Table 1: Descriptive Statistics — Job Market Candidates
N Mean Min Max
Gender
Female 603 0.255 0 1
Citizenship Region
Developing Countries 603 0.093 0 1
Former Soviet Union 603 0.217 0 1
North America 603 0.036 0 1
Europe 603 0.347 0 1
Asia 603 0.307 0 1
PhD Region
USA 603 0.498 0 1
Europe 603 0.323 0 1
Other English Speaking 603 0.164 0 1
Other 603 0.015 0 1
Research Field
Finance 603 0.272 0 1
Applied 603 0.342 0 1
Macroeconomics 603 0.232 0 1
Theory 603 0.100 0 1
Econometrics 603 0.055 0 1
PhD Ranking
1–25 603 0.252 0 1
26–50 603 0.235 0 1
51–75 603 0.083 0 1
76–100 603 0.075 0 1
101–150 603 0.141 0 1
151–200 603 0.081 0 1
201+ 603 0.133 0 1
PhD Duration
Seven years or more 603 0.096 0 1
Publications
At Least One Top Field 603 0.053 0 1
Notes: “Other English Speaking” includes UK, Ireland Rep., Australia, and Singapore. The PhD ranking is from Tilburg. I considered publications in JIE, JET, JoE, JME, JPubE, JEH, JFE, JF, Rand, JEEA, REStat, EJ, IER, all the AEJs, and the top 5.
Table 1 summarizes the main characteristics of the job market candidates in my data. There are some notable differences between my data and similar samples collected by Eberhardt et al. (2023) and Baltrunaite et al. (2024) from British and Italian institutions, respectively. Given that my institution is based in Russia, it is not surprising to see an over-representation of applicants from the former Soviet Union. [^26] Of these, 108 are Russian, 8 are Ukrainian, 5 are Armenian, 5 from the Baltic Countries, and 5 from other countries. There is also a relatively small fraction of North American citizens. About one in four applicants in my data are female, compared to around 30% in their samples. Another important difference is that the samples in their studies include both new PhDs and experienced academics. Because of this, the candidates in their samples have more publications.
In other respects, the samples are comparable. In particular, the fraction of candidates getting their PhD from an American institutions (around 50%) and the fraction in top-ranked programs are very similar.
3.2 Reference Letter Polarity
Figure 1 shows the distribution of polarity scores and letter length for four different dictionary approaches: the Harvard General Inquirer, the Loughran-McDonald lexicon, Vader, and Senticnet. Several facts are apparent. First, there is no clear relationship between letter length and letter sentiment. A locally weighted regression line is practically flat in all cases.
Second, advisers’ letters are distinctly longer than other referees’ letters. On average, non-adviser letters are in the neighborhood of 900 words long. Advisers’ letters are almost 33% longer on average, and the difference is highly statistically significant. Finally, note how several of the lexicons find no sentiment-carrying words in a sizeable number of cases. These cases appear as a horizontal line of zero-polarity for HGI, Senticnet and Vader.
Figure 1: The Distribution of Polarity and Letter Length — Dictionary Approaches
Notes: Fitted line is a locally weighted regression. LM polarity re-scaled.
In figure 2, I present similar scatter plots using prompt-based sentiment extraction for different variants of the Big Bird Model. The rightmost plot depicts the result of using the default sparse attention, which is the main innovation of the Big Bird model (see Zaheer et al., 2020, for details). The great advantage of sparse attention is that the context window is expanded to 4,096 tokens, which is enough to fit all the reference letters in my data. Unfortunately, as evident in the figure, sparse attention introduces a strong negative dependence between letter length and sentiment polarity. The model is simply not able to cope with the amount of information present in the long letters.
Big Bird can also be used in full attention mode, which limits the context window to 1,024 tokens. In order to process longer letters, I applied a technique known as windowing. In a nutshell, windowing splits a long document into partially overlapping “windows”. After some experimentation, I came to the conclusion that setting the stride —i.e. overlap— to half the window size (512 tokens) worked well. [^27] I also fine tuned the algorithm so that text windows contained only full sentences.
Figure 2: The Distribution of Polarity and Letter Length — Big Bird Prompting
Notes: Fitted line is a locally weighted regression.
Table 2 shows the results of applying the windowing technique. Most letters can be covered with up to two windows but a few require up to seven windows. Interestingly, the results suggest that the highest sentiment content in letters of reference is located at the end of the document. Frequent reference letter readers will not be surprised. The only remaining issue is how to summarize the up to seven scores per letter. As shown in the middle plot of figure 2, simply averaging the scores predictably biases the sentiment polarity of longer letters downwards. Therefore, I use the maximum score over all windows —invariably, this corresponds to the last window— as my final measure of letter sentiment.
Table 2: Windowing Results
Total Window Average Polarity $N$
Windows w1 w2 w3 w4 w5 w6 w7
1 0.256 835
2 0.064 0.215 560
3 0.054 0.035 0.268 281
4 0.064 0.024 0.052 0.290 133
5 0.091 0.019 0.021 0.037 0.210 44
6 0.092 0.011 0.013 0.030 0.075 0.259 29
7 0.171 0.007 0.021 0.037 0.024 0.082 0.315 16
Note: Prompt-based polarity scores.
In table 3, I present some descriptive statistics for the different sentiment measures. Reference letter sentiment is overwhelmingly positive, regardless of how one measures it. However, there is sizeable variation in exactly how positive the referee’s opinion is. Interestingly, half of the variance is explained by within-candidate variation — different referees having differing opinions on the same candidate.
Table 3: Sentiment Polarity Descriptives
A. Descriptive Stats
Big Bird Prompt Vader LM Senticnet HGI
$N$ 1898 1898 1898 1898 1898
$mean$ 0.256 0.245 0.539 0.310 0.476
$stddev$ 0.196 0.116 0.144 0.118 0.199
$within$ 0.636 0.536 0.441 0.521 0.499
$min$0.015 -0.166 0.000 0.000 -0.089
$max$ 0.847 0.780 1.000 0.630 1.000
B. Correlations
Big Bird Prompt 1.000 0.154 0.145 0.164 0.117
Vader 1.000 0.434 0.408 0.670
LM 1.000 0.430 0.531
Senticnet 1.000 0.485
Notes: Loughran-McDonald (LM) polarity is re-scaled. “Within” is the fraction of within individual variance.
Panel B of the table presents Pearson correlations among the different sentiment measures. While all correlations are positive, it is noteworthy that the lexicon-based measures have much stronger associations among each other —correlations close to 0.5— than with the prompt-based polarity score.
Figure 3: Prompt-based Sentiment Polarity Distribution
Notes: Horizontal lines mark the median score. Red crosses mark the mean.
Figure 3 shows how the distribution of prompt-based sentiment scores varies by different covariates. [^28] See appendix table A.1 for detailed descriptives for both sentiment scores and letter length. The job market is known to be quite dynamic. However, there is no evidence of a trend over time in average letter sentiment or its dispersion. There does seem to be an upwards trend in letter length, however.
I find a relatively small statistically significant difference in average sentiment scores ($\sim 3$ pp.) in favor of women. [^29] In contrast, the average length of letters written for male candidates is a bit higher ($\sim 20$ words); but the difference is not statistically significant.
A somewhat surprising finding involves differences across fields of research. Specifically, the letters written for candidates who work in theoretical fields (micro, game theory, mathematical economics) describe them in significantly less favorable terms than candidates in other fields. These kind of differences can have multiple explanations. Of course, one distinct possibility is that the theory applicants were on average a bit less competitive than applicants from other fields. Alternatively, it could be the case that for some reason referees are tougher on theorists. Finally, there could be something in my measure of sentiment that introduces a downward bias in theorists’ letters. [^30] For example, the fact that these letters tend to contain more mathematical expressions might make it harder for the LLM to process them. On the other hand, the letters for econometricians —which do not lack technicalities— score much better..
I do not find statistically significant differences across PhD program rank or region. Finally, candidates with North American or European citizenship tend to be scored a bit higher (anova p-value=0.056).
3.3 Testing Prompt-based Learning
Table 4: A Test of Prompt-based Learning — Predicting Candidate Sex
Predictive Performance
Precision Recall F1-score Support
Female 0.95 0.98 0.97 479
Male 0.99 0.98 0.99 1419
Accuracy 0.98 1898
Macro Avg. 0.97 0.98 0.98 1898
Weighted Avg. 0.98 0.98 0.98 1898
While it is possible to generate labels to evaluate a sentiment-scoring model, doing so is costly. Moreover, because of confidentiality and other considerations, it is particularly hard to label letters of recommendation.
Nevertheless, it is still possible to test the prompt-based procedure by predicting readily available features of the text. Here I provide an example consisting in predicting the sex of the job market candidate described in the letter.
[…]
Table 4 shows the predictive performance for this, admittedly simple, task. Basically, the model gets it right in all but a couple of cases.
4 Predicting Job Market Outcomes
There is no unambiguously appropriate measure of what constitutes “success” in the academic job market. Job market outcomes are heterogenous on multiple fronts. Consider figure 4, depicting different aspects of candidates’ placement by September after the job market season. Outcomes vary by the sector of employment, the type of position, and whether the job tasks are mainly research, teaching, or non-academic. In addition, employers differ on how they rank according to their research output, with higher ranked employers being markedly more desirable to the average junior academic.
Figure 4: Job Market Outcomes
Notes: Policy institutions include central banks, development agencies and think tanks. “Tenure Track” ={Assistant Prof., Lecturer, Senior Lecturer}.
I construct two outcome variables that are meant to encompass, as best as possible, the most desirable positions. I refer to the first outcome variable as “permanent research positions”. They include tenure-track research positions in academia and permanent research positions in policy institutions. The second outcome is the subset of permanent research positions in top-200 ranked universities and top-50 policy institutions. I refer to the latter as “elite positions”.
Figure 5: Dependent Variables
Notes: “Elite” positions are in the top-200 universities and top-50 policy institutions.
4.1 Linear Regression
In order to study the relation between a candidate’s reference letters and their job market outcome, I first look at the regression:
$y_{it}=$ $\>\alpha+\beta_{1}\,\mathrm{avg\_length}_{it}+\beta_{2}\,\mathrm{avg\_sentiment}_{it}+\boldsymbol{X}^{\prime}_{it}\boldsymbol{\gamma}+\lambda_{t}+\epsilon_{it}$
where $y_{it}$ is the job market outcome of candidate $i$, in job market year $t$. The main predictors in these models are the average letter length and average letter polarity. In all regressions, I introduce controls for the candidates’ characteristics (gender, nationality, publications, long PhD), their PhD program’s characteristics (rank, region) and the job market year.
Reference letters are only one input in the recruitment process. Job market outcomes also depend on the candidate’s ability favorably to present themselves in interviews, seminar presentations, and one-on-one meetings with faculty and administrators at potential employers, among other variables. Because these other determinants remain unobserved, the regression estimates must be interpreted with care. Assuming strong candidates get good letters and also tend to perform well in interviews and seminars, one reasonable starting point is to expect regression coefficients to over-estimate the true causal effect that the letters might have on a candidate’s prospects.
Table 5: Dictionary Methods — Predicting Job Market Outcomes
(1) (2) (3) (4)
HGI LM Vader Senticnet
A. Permanent Research Position
Average Letter Length 0.1658∗∗∗ 0.1600∗∗∗ 0.1653∗∗∗ 0.1640∗∗∗
(0.0531) (0.0531) (0.0532) (0.0544)
Average Polarity -0.0251 -0.1110 0.1141 -0.1853
(0.1298) (0.0855) (0.2298) (0.3064)
B. Elite Position
Average Letter Length 0.2780∗∗∗ 0.2771∗∗∗ 0.2766∗∗∗ 0.2807∗∗∗
(0.0513) (0.0515) (0.0514) (0.0521)
Average Letter Polarity 0.0720 -0.0145 0.2628 0.2349
(0.1222) (0.0788) (0.2121) (0.2385)
Candidate’s characteristics Yes
PhD characteristics Yes
Job Market Year Yes
Notes: $N=603$. Robust standard errors in parenthesis. p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01. Candidate’s characteristics: sex, nationality, publications, PhD lasting 7 years or longer. PhD program characteristics: rank category, region.
Table 5 presents a summary of my findings when using dictionary methods to measure letter sentiment. Across the board, I find that the lexical method produces sentiment scores that are not predictive of job market outcomes. Notice how, in contrast, average letter length has a strong positive association with the probability of successful placement in the job market. Letters that are on average about 1,000 words longer are associated with a 16–17 pp. higher probability of attaining a permanent research position, and about 28 pp. higher probability of employment in a high-ranked institution.
Table 6: Prompt-based Sentiment — Predicting Job Market Outcomes
Permanent Research Position Elite Position
Average Letter Length 0.1693∗∗∗ 0.2814∗∗∗
(000s words) (0.0524) (0.0509)
Average Prompt-based Polarity 0.5238∗∗∗ 0.5395∗∗∗
(0.1553) (0.1440)
Candidate’s characteristics Yes
PhD characteristics Yes
Job Market Year Yes
Notes: $N=603$. Robust standard errors in parenthesis. p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01. Candidate’s characteristics: sex, nationality, publications, PhD lasting 7 years or longer. PhD program characteristics: rank category, region.
In table 6, I present similar regression estimates when using my prompt-based strategy. The results for average letter length are almost identical, which is not too surprising given that letter length and letter polarity are not strongly correlated. On the other hand, in contrast to dictionary methods, the prompt-based sentiment is strongly predictive of good job market outcomes.
One point worth emphasizing is how prompt-based learning enables a straightforward interpretation of the regression results, even if they cannot be taken to measure a causal effect. For example, an increase in average polarity of 0.1 means that the LLM assigns a 10 pp. higher probability to filling the prompt with one of the positive sentiment tokens relative to the negative tokens. According to the findings in the table, such as increase is predictive of a 5.2 pp. increase in the expected probability of obtaining a permanent research position, and a 5.4 pp. higher probability of obtaining an “elite” position.
4.2 Nonlinearities
In order to explore nonlinear relationships and interactions, I estimate random forest classifiers with the same outcome and predictor variables. [^31] See Athey and Imbens (2019) for details on random forests. The main hyper-parameters are set based on 5-fold cross-validation. [^32] Table A.2 in the appendix presents the hyper-parameter values. As usual, the estimation of each decision tree is based on a different sample bootstrap and a random subset of the predictors.
I use the fitted random forests to obtain partial dependence curves, which simulate the changes in the predicted probability for each of the outcomes. Formally, the partial dependence curve for the predictors $X_{S}$ is obtained as follows:
$\mathrm{pd}_{X_{S}}(x_{S})=\mathrm{E}_{X_{C}}\left[\hat{P}(x_{S},X_{C})\right],\quad x_{S}\in\mathcal{X}_{S}$
where $X_{C}$ are all other predictors and $\mathcal{X}_{S}$ is the support set of $X_{S}$.
In figure 6, I present partial dependence curves for the average letter prompt-based sentiment score. While the random forest confirms the positive association between letter scores and employment prospects, it suggests that the relationship might not be linear.
Figure 6: Partial Dependence Curves — Average Prompt-based Letter Polarity
(a) Permanent Research Position
(b) Elite Position
Note: Partial dependence curves simulate average job market outcomes when the whole sample is assigned (counterfactual) levels of average letter polarity.
The same method can be applied to study interactions between letter sentiment and other predictors. For example, to study heterogeneity across gender lines, I let $X_{S}=\{\mathrm{avg\_sentiment},\mathrm{gender}\}$. The results of these simulations are presented in figure 7.
Figure 7: Interactions — Letter Polarity and Gender
(a) Permanent Research Position
(b) Elite Position
Note: Partial dependence curves simulate average job market outcomes when the whole sample is assigned (counterfactual) values for gender and average letter polarity.
The average predicted probability of obtaining a permanent research position is virtually identical when the sample is assigned either sex, regardless of the letter sentiment. In contrast, women are predicted to have a small advantage in getting hired at top institutions, with the sex gap somewhat narrowing as letters get stronger.
Figure 8: Interactions — Letter Polarity and Field
(a) Permanent Research Position
(b) Elite Position
Note: Partial dependence curves simulate average job market outcomes when the whole sample is assigned (counterfactual) values for field of research and average letter polarity.
Figures 8 and 9 present similar simulations, this time by field of research and PhD rank group respectively.
Figure 9: Interactions — Letter Polarity and PhD Rank
(a) Permanent Research Position
(b) Elite Position
Note: Partial dependence curves simulate average job market outcomes when the whole sample is assigned (counterfactual) values for PhD rank group and average letter polarity.
Clearly, applied fields have a tougher time in the market. Even candidates with very strong letters have distinctly lower predicted probability of attaining a permanent research position or an elite position. In fact, the gap with other fields seems to widen as letters get stronger.
Interestingly, the random forest predictions suggest that the ranking of the PhD program does not make a very sizeable difference in terms of obtaining a research position. On the other hand, elite positions are easier to access for PhD candidates from top ranked universities. The gradient with respect to average letter polarity is similar for all candidates regardless of their PhD rank.
Figure 10: Interactions — Letter Polarity and Letter Length
(a) Permanent Research Position
(b) Elite Position
Note: Partial dependence contours depict average job market outcomes when the whole sample is assigned (counterfactual) levels of average letter polarity and length.
A final point involves interaction between letter sentiment and letter length. Figure 10 shows that, predictably, these two forms of signalling positively interact
5 Preliminary Conclusions
There is compelling evidence that reference letters are predictive of job market outcomes. Both average letter length and sentiment polarity are positively associated with the probability of success in general research positions. The relationship seems to be even stronger for more selective “elite” positions.
All in all, the evidence seems to be consistent with the canonical Saloner (1985) model. There is competition among referees to place their students in good places, and this creates incentives to somewhat “inflate” candidates’ achievements and prospects. However, the repetitive nature of the interaction between letter-writers and employers, plus the fact that individual ability is eventually revealed mean that there are counterbalancing forces. The resulting equilibrium is one in which relatively stronger candidates get better recommendations, employers place significant trust in the contents of the references, and those candidates are more likely to get the best job offers.
This meritocratic mechanism seems to operate quite broadly, across fields of research, personal characteristics of the candidate and throughout the rankings. In fact, one of the most refreshing takeaways from the findings is that good candidates from relatively lower ranked PhDs can offset a sizeable part of any starting point disadvantage by producing research worthy of excellent referee reports.
On the methodological front, a key takeaway of the evidence presented here is that dictionary methods have limited power. Because they do not capture context, they are unable to extract some relevant information. The prompt-based strategy described here is a simple enough alternative that leverages the enormous resources invested in pre-training LLMs. Extending and refining this method should be the focus of further research.
Appendix A Appendix
Table A.1: Sub-group Sentiment Polarity and Letter Length
BB Prompt-based Sentiment (pp.) Word Count (,000s)
Period
2013 0.262 0.845
2015 0.250 0.949
2016 0.236 0.880
2017 0.229 0.977
2018 0.253 0.924
2019 0.259 1.060
2020 0.265 1.058
2021 0.280 1.197
p-val 0.164 0.000
Gender
Female 0.277 0.982
Male 0.249 1.003
p-val 0.006 0.458
Field
Applied 0.261 1.042
Econometrics 0.235 0.959
Finance 0.264 0.896
Macroeconomics 0.279 1.000
Theory 0.174 1.144
p-val 0.000 0.000
PhD Rank
1–25 0.238 1.077
26–50 0.255 1.036
51–75 0.257 0.937
76–100 0.255 0.972
101–150 0.253 0.988
151–200 0.277 1.003
201+ 0.283 0.836
p-val 0.109 0.000
Citizenship Region
Asia 0.247 0.948
Developing 0.238 0.998
Europe 0.270 1.044
Former Soviet 0.248 0.995
North America 0.296 1.007
p-val 0.056 0.045
PhD Region
USA 0.253 0.989
English Speaking 0.275 0.991
Europe 0.252 1.021
Other 0.252 0.864
p-val 0.296 0.376
Notes: The p-val correspond to an anova test of equality of means across groups.
Table A.2: Random Forest Hyper-Parameters
Permanent Research Position Elite Position
Max Depth 4 5
Min Samples to Split Node 30 29
Min Samples in Leaf Node 13 9
Forest Size 140 150
Notes: Hyper-parameters result from 5-fold CV search. Samples are bootstrapped and $\sqrt{9}$ predictors are selected at random in the CART tree estimation. The impurity measure is the Gini.